OpenStudy (anonymous):

The time needed to travel a distance of y units if the speed x is reduced by 2.

OpenStudy (ahsome):

What is the relationship of x to y in the first place?

OpenStudy (anonymous):

S=D/T T=D/S = y / (x-2)

OpenStudy (ahsome):

^^ Thats it. You can't find the number unless more info is given

OpenStudy (anonymous):

true, just using the info given